OpenStudy (anonymous):

TRUE/FALSE The value 5pi/2 is a solution for the equation...

OpenStudy (anonymous):

\[2\sin ^{2}x-\sin x -1 =0\]

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

did u try plugging in 5pi/2 ? you can allso solve for sin x if you want... plug in sin x = y and u get a quadratic in y

OpenStudy (anonymous):

Is this right? \[2\sin ^{2}(90)-\sin(90)-1=0?\] @hartnn

hartnn (hartnn):

90 ?? 5pi/2 isn't 90 .....

hartnn (hartnn):

ok, its equivalent to 90, you can use 90 instead

hartnn (hartnn):

so, whats sin 90 =... ?

OpenStudy (anonymous):

sin(90)=1

hartnn (hartnn):

so, 2 (1) -1 -1 does this = 0 ?

OpenStudy (anonymous):

Yes!!

hartnn (hartnn):

then 5pi/2 is the solution :)

OpenStudy (anonymous):

Thanks!!

hartnn (hartnn):

welcomes !!